# Parallel-Probe: Towards Efficient Parallel Thinking via 2D Probing

### v1

@@ 1 Introduction
To improve efficiency, previous work studies efficient reasoning at test time. The majority of the research investigates early-stopping strategies for sequential generation (e.g., extended Chain-of-Thought), leveraging signals such as confidence (Fu et al., 2025b), hidden states (Li et al., 2026), or answer convergence (Liu and Wang, 2025; Zhang et al., 2025b). Since these approaches focus on the internal state of individual trajectories, they ignore critical global information across branches (e.g., consensus), making them sub-optimal in parallel thinking settings. Meanwhile, several studies have explored adaptive sampling to reduce the inference cost of self-consistency (Mao et al., 2025; Aggarwal et al., 2023; Wan et al., 2025; Fu et al., 2025b; Huang et al., 2025). Since these methods rely on sequential control loops, they transform parallel sampling into a semi-sequential process. Consequently, even though sample efficiency is improved, the increased latency cancels out the speed advantage. Efficient parallel thinking in an online setting has received limited attention, particularly the simultaneous launch of multiple paths.
@@ 2.2 Observations From 2D Probing
These observations expose a fundamental mismatch in current parallel thinking designs: while reasoning branches are executed independently, the “signal” is a collective, global property. Traditional stopping criteria, which rely on local trajectory signals (e.g., confidence (Fu et al., 2025b) or answer convergency (Zhang et al., 2025b; Liu and Wang, 2025)), fail to capture this cross-branch consensus. Consequently, a new set of principles are required to shift control from individual trajectories to the global dynamics of the parallel thinking.
@@ 5.3 Baseline Methods
  • SC@64 + SAC (Liu and Wang, 2025): A baseline that applies SAC as a trajectory-level early stopping rule within SC, terminating each trajectory upon local answer convergence before majority voting.
@@ 7.1 Efficient Parallel Reasoning
To mitigate the computational cost of fixed-budget search, recent research focuses on dynamic resource allocation. Aggarwal et al. (2023) and Li et al. (2024) propose adaptive mechanisms that halt generation once a consensus threshold is met, while Wang et al. (2025b) further optimizes efficiency by allocating samples based on query difficulty. Beyond count reduction, confidence-aware approaches weight reasoning paths to identify high-quality solutions with fewer samples (Huang et al., 2025; Taubenfeld et al., 2025; Fu et al., 2025b). However, they predominantly adopt sequential sampling to obtain these samples, limiting the hardware efficiency of parallel thinking. More recently, fine-grained methods like Dynamic Self-Consistency (Wan et al., 2025), Self-Truncation (Wang et al., 2025c), DeepPrune (Tu et al., 2025), Step (Liang et al., 2026) and Slim-SC (Hong et al., 2025) prune unpromising trajectories mid-generation to minimize wasteful computation on incorrect paths. Despite their effectiveness, these methods lack principled modeling of the global dynamics across parallel reasoning trajectories, resulting in coarse-grained control over parallel thinking.
@@ 7.2 Efficient Sequential Reasoning
To optimize the depth of thought without additional training, recent research focuses on dynamic early exiting mechanisms. A primary strategy involves monitoring uncertainty metrics: Wang et al. (2025a) and Sharma and Chopra (2025) utilize entropy after the reasoning block or at the sequence level as confidence signals, while Yong et al. (2025) estimates this empirically via multiple rollouts or beam search. Alternatively, termination decisions can be guided by output stability, using answer convergence across steps to signal sufficiency (Liu and Wang, 2025; Mao et al., 2025; Fu et al., 2025a; Zhang et al., 2025b). Beyond output statistics, Zhang et al. (2025a) suggest probing hidden states directly for self-verification, allowing models to halt inference once an internal correctness threshold is met (Yang et al., 2025b). Despite their success in efficient sequential reasoning, these methods fail to leverage the global dynamics of parallel thinking (as reflected in our Observations 1–3). As a result, directly applying them to parallel reasoning settings is sub-optimal.
@@ 7.3 Test-Time Scaling
To optimize the efficiency of complex reasoning, recent studies have shifted focus toward the strategic allocation of test-time computation (Snell et al., 2024; Chen et al., 2025b; Xiong et al., 2025). A primary manifestation of this trend is the use of tree-search frameworks, which aggregate diverse reasoning paths and employ sparse activation to manage complexity (Bi and others, 2024; Lample et al., 2022; Koh et al., 2024; Zheng et al., 2025). To further refine these search spaces, step-wise verifiers have become essential for dynamically pruning unproductive branches (Wang et al., 2022b; Li et al., 2022; Lightman et al., 2023). Beyond search-level optimizations, performance can be bolstered by diversifying query formulations (Huang et al., 2024) or through iterative refinement cycles that bootstrap the model’s self-correction capabilities to handle increasingly intricate tasks (Chen et al., 2025a; Welleck et al., 2022; Madaan et al., 2023; Aggarwal et al., 2024). Our work leverages global dynamic signals from black-box 2D probing to enable principled control along both depth and width dimensions.

### v2

@@ 1 Introduction
To improve efficiency, previous work studies efficient reasoning at test time. The majority of the research investigates early-stopping strategies for sequential generation (e.g., extended Chain-of-Thought), leveraging signals such as confidence (Fu et al., 2025b), hidden states (Li et al., 2026), or answer convergence (Liu & Wang, 2025; Zhang et al., 2025b). Since these approaches focus on the internal state of individual trajectories, they ignore critical global information across branches (e.g., consensus), making them sub-optimal in parallel thinking settings. Meanwhile, several studies have explored adaptive sampling to reduce the inference cost of self-consistency (Mao et al., 2025; Aggarwal et al., 2023; Wan et al., 2025; Fu et al., 2025b; Huang et al., 2025). Since these methods rely on sequential control loops, they transform parallel sampling into a semi-sequential process. Consequently, even though sample efficiency is improved, the increased latency cancels out the speed advantage. Efficient parallel thinking in an online setting has received limited attention, particularly the simultaneous launch of multiple paths.
@@ 2.2 Observations From 2D Probing
These observations expose a fundamental mismatch in current parallel thinking designs: while reasoning branches are executed independently, the “signal” is a collective, global property. Traditional stopping criteria, which rely on local trajectory signals (e.g., confidence (Fu et al., 2025b) or answer convergency (Zhang et al., 2025b; Liu & Wang, 2025)), fail to capture this cross-branch consensus. Consequently, a new set of principles are required to shift control from individual trajectories to the global dynamics of the parallel thinking.
@@ 5.3 Baseline Methods
  • SC@64 + SAC (Liu & Wang, 2025): A baseline that applies SAC as a trajectory-level early stopping rule within SC, terminating each trajectory upon local answer convergence before majority voting.
@@ 7.1 Efficient Parallel Reasoning
To mitigate the computational cost of fixed-budget search, recent research focuses on dynamic resource allocation. Aggarwal et al. (2023) and Li et al. (2024) propose adaptive mechanisms that halt generation once a consensus threshold is met, while Wang et al. (2025a) further optimizes efficiency by allocating samples based on query difficulty. Beyond count reduction, confidence-aware approaches weight reasoning paths to identify high-quality solutions with fewer samples (Huang et al., 2025; Taubenfeld et al., 2025; Fu et al., 2025b). However, they predominantly adopt sequential sampling to obtain these samples, limiting the hardware efficiency of parallel thinking. More recently, fine-grained methods like Dynamic Self-Consistency (Wan et al., 2025), Self-Truncation (Wang et al., 2025c), DeepPrune (Tu et al., 2025), Step (Liang et al., 2026) and Slim-SC (Hong et al., 2025) prune unpromising trajectories mid-generation to minimize wasteful computation on incorrect paths. Despite their effectiveness, these methods lack principled modeling of the global dynamics across parallel reasoning trajectories, resulting in coarse-grained control over parallel thinking.
@@ 7.2 Efficient Sequential Reasoning
To optimize the depth of thought without additional training, recent research focuses on dynamic early exiting mechanisms. A primary strategy involves monitoring uncertainty metrics: Wang et al. (2025b) and Sharma & Chopra (2025) utilize entropy after the reasoning block or at the sequence level as confidence signals, while Yong et al. (2025) estimates this empirically via multiple rollouts or beam search. Alternatively, termination decisions can be guided by output stability, using answer convergence across steps to signal sufficiency (Liu & Wang, 2025; Mao et al., 2025; Fu et al., 2025a; Zhang et al., 2025b). Beyond output statistics, Zhang et al. (2025a) suggest probing hidden states directly for self-verification, allowing models to halt inference once an internal correctness threshold is met (Yang et al., 2025b). Despite their success in efficient sequential reasoning, these methods fail to leverage the global dynamics of parallel thinking (as reflected in our Observations 1–3). As a result, directly applying them to parallel reasoning settings is sub-optimal.
@@ 7.3 Test-Time Scaling
To optimize the efficiency of complex reasoning, recent studies have shifted focus toward the strategic allocation of test-time computation (Snell et al., 2024; Chen et al., 2025b; Xiong et al., 2025). A primary manifestation of this trend is the use of tree-search frameworks, which aggregate diverse reasoning paths and employ sparse activation to manage complexity (Bi et al., 2024; Lample et al., 2022; Koh et al., 2024; Zheng et al., 2025). To further refine these search spaces, step-wise verifiers have become essential for dynamically pruning unproductive branches (Wang et al., 2022b; Li et al., 2022; Lightman et al., 2023). Beyond search-level optimizations, performance can be bolstered by diversifying query formulations (Huang et al., 2024) or through iterative refinement cycles that bootstrap the model’s self-correction capabilities to handle increasingly intricate tasks (Chen et al., 2025a; Welleck et al., 2022; Madaan et al., 2023; Aggarwal et al., 2024). Our work leverages global dynamic signals from black-box 2D probing to enable principled control along both depth and width dimensions.